OpenStudy (anonymous):

i am in algebra and im having trouble with this problem it is 3x=-2y-5 2y=-5x+5 i am using substitution method

OpenStudy (anonymous):

Ok so because you are using substitution you want to get either y or x by itself. Which one would you like to get by itself?

OpenStudy (anonymous):

X

OpenStudy (anonymous):

Alright so... 3x=-2y-5 divide each side by 3 to get x alone x=-2/3y-5/3

OpenStudy (anonymous):

Now you can take this and plug it into your other equation to substitute for x. 2y=-5x+5 2y=-5(-2/3y-5/3 )+5 2y=10/3y+40/3 combine like terms (so subtract 10/3y from both sides) -4/3y=40/3 divide each side by -4/3 y=-10

OpenStudy (anonymous):

Now that you know your y value take y, which is -10 and plug it into your original equation 3x=-2y-5 3x=-2(-10)-5 3x=20-5 3x=15 divide each side by 3 x=5

OpenStudy (anonymous):

y=-10 x=5

OpenStudy (anonymous):

thank you Buckky

OpenStudy (anonymous):

No problem hope it makes sense